OpenStudy (anonymous):

r^2+13r-30=0 step by step please with final answer thanks

jimthompson5910 (jim_thompson5910):

r^2+13r-30=0 (r+15)(r-2)=0 r+15=0 or r-2=0 r=-15 or r=2

OpenStudy (anonymous):

okay can you check it for me please, thanks

jimthompson5910 (jim_thompson5910):

Simply plug r=-15 into the original equation and simplify. You should get 0=0 as a result. Do the same for r=2.

OpenStudy (anonymous):

thanks